A 25.0L tank contains 2.0kg of neon gas with a pressure of 98.0 atm at 28 degrees celcius. How many kilograms of oxygen can fill the tank under the same pressure and temperature?

1
Expert's answer
2021-11-11T11:58:41-0500

Solution.

"V=25.0\sdot 10^{-3}m^3;"

"m=2.0 kg;"

"M(Ne)=20.2\sdot 10^{-3}kg/mol;"

"M(O_2)=32\sdot10^{-3} kg/mol;"

"p=98.0 atm=9929850 Pa;"

"T=301K;"

"pV=\dfrac{m(Ne)}{M(Ne)}RT\implies V=\dfrac{m(Ne)}{M(Ne)p}RT;"

"V=\dfrac{2.0}{20.2\sdot10^{-3}\sdot9929850}\sdot8.31\sdot301=0.025 m^{-3};"

"p=const; T=const;"

"pV=\dfrac{m(O_2)}{M(O_2)}RT\implies m(O_2)=\dfrac{pVM(O_2)}{RT};"

"m(O_2)=\dfrac{9929850\sdot25.0\sdot10^{-3}\sdot32\sdot10^{-3}}{8.31\sdot301}=3.2kg;"


Answer: "m(O_2)=3.2kg."